### v1

# Interlayer interactions in $\text{La}_{3}\text{Ni}_{2}\text{O}_{7}$ under pressure: from $s^{\pm}$ to $d_{xy}$-wave superconductivity

Abstract
We investigate the role of interlayer interaction terms in the competition between different superconducting gap symmetries in the bilayer nickelate $\text{La}_{3}\text{Ni}_{2}\text{O}_{7}$ under high pressure. We study a two-layer, two-orbital electron model that encompasses both intra- and interlayer Coulomb interaction terms within the matrix random-phase approximation. We find that interlayer interactions favor a $d_{xy}$-wave superconducting pairing symmetry over the $s^{\pm}$-wave symmetry, which has been found to prevail when interlayer interactions are disregarded. Moreover, our findings indicate that interlayer interactions enhance the interorbital pairing, incorporating contributions from all three electron pockets, arising from both $d_{3z^{2}-r^{2}}$ and $d_{x^{2}-y^{2}}$ orbital character, resulting in nodes within the gap function (not present in the $s^{\pm}$-wave state) and consequently favoring the $d_{xy}$-wave pairing.
The recent discovery of high-temperature superconductivity in the Ruddlesden-Popper bilayer nickelate compound $\text{La}_{3}\text{Ni}_{2}\text{O}_{7}$ (LNO327) under high pressure [1, 2, 3] has sparked intense research on this new family of high-$T_{c}$ materials [4]. Despite the rapid progress in the last two years, including recent reports of superconductivity on thin films at ambient pressure [5], several fundamental questions remain on the nature of the underlying pairing mechanism for superconductivity in these compounds.
For example, there has been an intense debate as to which low-energy minimal model to describe the superconducting phase. Density Functional Theory calculations [6, 7, 8, 9] reveal that multiple bands, mainly formed by Ni-centered in-plane $d_{x^{2}-y^{2}}$ and out-of-plane $d_{3z^{2}-r^{2}}$ orbitals, along with oxygen-centered $p$-orbitals cross the Fermi level, forming two electron pockets ($\alpha$, $\beta$) with mixed orbital character and one hole pocket ($\gamma$) primarily associated with the Ni-$d_{3z^{2}-r^{2}}$ orbital [6]. These bands can be described by a two-layer, two-orbital minimal tight-binding model [6, 7, 9], which has been investigated with interlayer interactions within a localized electron picture [10] and without interlayer interactions in an itinerant electron picture [11, 12, 8, 9, 13]. In all cases, RPA calculations suggest that $s^{\pm}$-wave pairing is the dominant superconducting instability, driven by intra-orbital ($\pi$, 0) scattering between the $\beta$ and the $\gamma$-pockets [9, 8]. However, the d-wave pairing channel is found to have energy close to the s-wave pairing channel [12, 8].
Other works [14, 15] propose a minimal model containing just one of these orbitals (either $d_{x^{2}-y^{2}}$ or $d_{3z^{2}-r^{2}}$), which predict an $s^{\pm}$-wave superconducting gap symmetry, with or without interlayer interactions. In contrast, a proposal of cuprate-like $d$-$p$ hybridized single orbital brought up the possibility of $d_{x^{2}-y^{2}}$ or $d_{x^{2}-y^{2}}+is$ superconductivity [16].
On the experimental side, recent explorations of the ambient pressure properties of LNO327 give indications that interlayer interactions can be relevant for the superconducting phase at high pressures. Neutron scattering experiments in polycrystalline samples show evidence of spin excitations in the inelastic channel, which are consistent with strong interlayer exchange couplings [17]. In addition, X-ray absorption measurements [18] revealed dispersive magnons and spin-density wave order patterns which are also consistent with interlayer magnetic exchange being about an order of magnitude larger than intralayer ones. In this work, we investigate the role of the interlayer interaction terms in the competition between different superconducting gap symmetries in the system. We consider a two-layer, two-orbital electron model that includes both intra- and interlayer Coulomb interaction terms. Within the matrix random-phase approximation (mRPA) [19, 20, 21, 22, 23], we find that interlayer interactions favor a $d_{xy}$-wave superconductivity over the $s^{\pm}$ state, which is known to dominate when such interlayer interactions are neglected [11, 8, 14]. Our mRPA calculations reveal that interlayer interactions enhance interorbital pairing involving contributions from all three pockets, creating nodes in the gap function (absent in the $s^{\pm}$-wave state) and thus favoring the $d_{xy}$-wave pairing.
Band structure.— In order to describe the (non-interacting) band structure of the system, we consider the four-band low-energy model of Refs. [6, 12]. This tight-binding model involves two orbitals ($d_{x^{2}-y^{2}}$ and $d_{3z^{2}-r^{2}}$) per Ni atom in a unit cell with two Ni atoms (one for each layer) [as depicted in Fig. 1(a)] and hoppings up to second nearest neighbors. For a valence of 2.5 electrons per Ni atom [24], the resulting Fermi surface and band structure are shown in Fig. 1(b) and (c), respectively, together with the respective Ni orbital contributions.
The FS shows electron pockets $\alpha$ and $\beta$ with a strong contribution from the $d_{x^{2}-y^{2}}$ orbital (red), consistent with ARPES results at ambient pressure [25, 26]. The $\gamma$ hole pocket at $(\pi,\pi)$, however, is dominated by the $d_{3z^{2}-r^{2}}$ orbital (blue) and its presence has been linked to the onset of superconductivity if LNO327 and similar materials [8, 27, 28]. We notice that some of the bands show mixed orbital regions (nearly white) along the $\bm{\Gamma}-\bm{X}$ and $\bm{X}-\bm{M}$ high-symmetry paths [Fig. 1(c)].
Interacting model.— Our model Hamiltonian $H=H_{0}+H_{I}$ involves both the (non-interacting) part described by $H_{0}$ as well as an interacting term ($H_{I}$) that includes both onsite ($U,J$) and interlayer ($U_{\perp},J_{\perp}$) Coulomb and exchange interaction terms, as represented in Fig. 1(a).
The interacting Hamiltonian for the two-orbital itinerant electron model is given by the onsite plus interlayer parts, $H_{I}=H_{\text{onsite}}+H_{\text{inter}}$, where the onsite term is given by
$\begin{split}H_{\text{onsite}}&=U\sum_{ij,p\ell}n_{ip\ell\uparrow}n_{ip\ell\downarrow}\\ &+(U^{\prime}-J/2)\sum_{i\ell,p<q}n_{ip\ell}n_{iq\ell}\\ &-2J\sum_{i\ell,p<q}\bm{S}_{ip\ell}\cdot\bm{S}_{iq\ell}\\ &+J\sum_{i\ell,p<q}\left(c^{\dagger}_{ip\ell\uparrow}c^{\dagger}_{ip\ell\downarrow}c_{iq\ell\downarrow}c_{iq\ell\uparrow}+\text{h.c.}\right)\end{split}$ (1)
and the interlayer term reads
$\begin{split}H_{\text{inter}}&=U_{\perp}\sum_{ij,p\sigma\bar{\sigma}}n_{ipA\sigma}n_{ipB\bar{\sigma}}\\ &+(U_{\perp}^{\prime}-J_{\perp}/2)\sum_{i,p<q}n_{ipA}n_{iqB}\\ &-2J_{\perp}\sum_{i,p<q}\bm{S}_{ipA}\cdot\bm{S}_{iqB}\\ &+J_{\perp}\sum_{i,p<q}\left(c^{\dagger}_{ipA\uparrow}c^{\dagger}_{ipB\downarrow}c_{iqB\downarrow}c_{iqA\uparrow}+\text{h.c.}\right).\end{split}$ (2)
In the above, $c^{\dagger}_{ip\ell\sigma}$ creates an electron with spin $\sigma$, in orbital $p$, layer $\ell=A,B$, at the (square) lattice site $i$. In addition, we use a shortcut notation for the density operator, $n_{ip\ell}=\sum_{\sigma}n_{ip\ell\sigma}$. In the present model, $U(U_{\perp})$, $U^{\prime}(U_{\perp}^{\prime})$, and $J(J_{\perp})$ are the onsite (intersite) Hubbard, interorbital, and exchange interaction strengths, respectively.
Figure 1: Schematic representation of the bilayer $\text{La}_{3}\text{Ni}_{2}\text{O}_{7}$ model (a) with onsite $U,J$ and interlayer $U_{\perp},J_{\perp}$ Coulomb and exchange interaction strengths. Fermi surface (b) and band structure (c) for the two-orbital $d_{3z^{2}-r^{2}}$ (blue color) and $d_{x^{2}-y^{2}}$ (red color) low-energy model.
The interacting part of the Hamiltonian, $H_{I}$, can conveniently be rewritten in momentum space as
$\begin{split}H_{I}&=\sum_{\bm{q}}\sum_{pqst\ell\ell^{\prime}}(U_{c})^{p\ell q\ell}_{s\ell^{\prime}t\ell^{\prime}}(\bm{q})n_{p\ell q\ell}(-\bm{q})n_{s\ell^{\prime}t\ell^{\prime}}(\bm{q})\\ &+\sum_{\bm{q}}\sum_{pqst\ell\ell^{\prime}}(U_{s})^{p\ell q\ell}_{s\ell^{\prime}t\ell^{\prime}}(\bm{q})\bm{S}_{p\ell q\ell}(-\bm{q})\cdot\bm{S}_{s\ell^{\prime}t\ell^{\prime}}(\bm{q}),\end{split}$ (3)
where $n_{pAqA}(\bm{q})=\sum_{\bm{k}\sigma}c^{\dagger}_{(\bm{k}+\bm{q})pA\sigma}c_{\bm{k}qA\sigma}$ is the charge density and $\bm{S}_{pAqA}(\bm{q})=\sum_{\bm{k}\sigma\bar{\sigma}}c^{\dagger}_{(\bm{k}+\bm{q})pA\sigma}\bm{\sigma}_{\sigma\bar{\sigma}}c_{\bm{k}qA\bar{\sigma}}$ the spin density for layer $A$ (similarly for layer $B$), with $\bm{\sigma}_{\sigma\bar{\sigma}}$ being the Pauli vector. The matrix elements of the charge and spin interaction matrices which connect Eq. (3) with Eqs. (1) and (2) are given by
$\begin{split}&(U_{s})^{pApA}_{pApA}=U,\;(U_{s})^{pApA}_{qAqA}=U^{\prime},\\ &(U_{s})^{pAqA}_{pAqA}=(U_{s})^{pAqA}_{qApA}=J,\\ &(U_{c})^{pApA}_{pApA}=U,\;(U_{c})^{pApA}_{qAqA}=(-U^{\prime}+2J),\\ &(U_{c})^{pAqA}_{qApA}=J,\;(U_{c})^{pAqA}_{pAqA}=(-J+2U^{\prime}),\\ &(U_{c})^{pApA}_{pBpB}=2U_{\perp},\;(U_{c})^{pApA}_{qBqB}=2U_{\perp}^{\prime},\\ &(U_{c})^{pAqA}_{pBqB}=(U_{c})^{pAqA}_{qBpB}=2J_{\perp},\\ \end{split}$ (4)
where $A,B$ denote the respective layer index  [^1] The remaining matrix elements can be trivially obtained by exchanging $A\leftrightarrow B$. For simplicity, we assume spin rotational invariance such that the standard relationship $U^{\prime}=U-2J$ holds and we use the usual relation $J=U/4$ and $J_{\perp}=U_{\perp}/4$. Notice the interlayer interactions only play a role on the charge interaction channel.
We now consider the singlet-channel RPA pairing interaction matrix [19, 21] for the multi-orbital, multilayer interacting model of Eqs. (3) and (4) [30]:
$\begin{split}\Gamma^{p\ell q\ell^{\prime}}_{s\xi t\xi^{\prime}}(\bm{q})=\frac{1}{2}&\Big{[}-\hat{U}_{c}\hat{\chi}_{c}(\bm{q})\hat{U}_{c}\\ &+3\hat{U}_{s}\hat{\chi}_{s}(\bm{q})\hat{U}_{s}+\hat{U}_{c}+\hat{U}_{s}\Big{]}^{t\xi^{\prime}q\ell^{\prime}}_{p\ell s\xi}\end{split}$ (5)
where $\hat{\chi}_{s}$ and $\hat{\chi}_{c}$ are the static spin and charge susceptibility matrices, respectively.
Within the mRPA approximation, the pairing interaction $\Gamma(\bm{k},\bm{k}^{\prime})$ between Cooper pairs at $(\bm{k},-\bm{k})$ and $(\bm{k}^{\prime},-\bm{k}^{\prime})$ can be computed from Eq. (5) as [19, 20]
$\begin{split}\Gamma(\bm{k},\bm{k}^{\prime})=\text{Re}&\sum_{\ell\ell^{\prime}\xi\xi^{\prime}}\sum_{pqst}\left[\psi^{t\xi^{\prime}}_{\nu_{-\bm{k}}}(-\bm{k})\right]^{*}\left[\psi^{s\xi*}_{\nu_{\bm{k}}}(\bm{k})\right]^{*}\\ &\times\Gamma^{p\ell q\ell^{\prime}}_{s\xi t\xi^{\prime}}(\bm{k}-\bm{k}^{\prime})\psi^{p\ell}_{\nu_{\bm{k}^{\prime}}}(\bm{k}^{\prime})\psi^{q\ell^{\prime}}_{\nu_{-\bm{k}^{\prime}}}(-\bm{k}^{\prime})\end{split}$ (6)
where $\psi^{p\ell}_{\nu_{\bm{k}}}(\bm{k})$ is the $\nu_{\bm{k}}$-th band single-particle electronic state with momentum $\bm{k}$ in a basis where $p$ and $\ell$ are good quantum numbers. The symmetric part of the pairing interaction $(\bar{\Gamma}(\bm{k},\bm{k}^{\prime}))$ can, in turn, be used to calculate the gap function $\Delta^{\alpha}(\bm{k})$ and the respective pairing strength $\lambda^{\alpha}$ using the usual integral eigenvalue equation [19]:
$-\sum_{j}\oint_{\bm{k^{\prime}}\in C_{j}}\frac{d\bm{k}^{\prime}_{||}}{v_{F}(\bm{k}^{\prime})}\frac{1}{(2\pi)^{2}}\bar{\Gamma}(\bm{k},\bm{k}^{\prime})\Delta^{\alpha}(\bm{k^{\prime}})=\lambda^{\alpha}\Delta^{\alpha}(\bm{k})\;,$ (7)
where the sum runs over the different FS sheets $C_{j}$ and $v_{F}(\bm{k})$ denotes the Fermi speed at momentum $\bm{k}$.
Figure 2: (a) Phase diagram showing the log of the ratio between the two leading eigenvalues of Eq. (7) for given values of $U$ and $U_{\perp}$. The corresponding leading gap functions $\Delta_{\rm max}(\bm{k})$ displaying $s^{\pm}$-wave and $d_{xy}$-wave symmetries for $U=1.16$ eV and $U_{\perp}/U=0.1$ (triangle) and $U_{\perp}/U=0.3$ (square) are shown in panels (b) and (c), respectively.
Superconducting symmetries.— For a given set of parameters entering $\bar{\Gamma}(\bm{k},\bm{k}^{\prime})$, the solution of Eq. (7) with the highest eigenvalue $\lambda_{\rm max}$ will correspond to the highest critical temperature $T_{c}$ and the symmetry of the pairing gap is reflected in the corresponding eigenfunction $\Delta_{\rm max}(\bm{k})$. In the following, we consider the reference value of $U\approx 1.16$ eV, which has been estimated to match $T_{c}\approx 80$ K [12]. Considering the largest interlayer hopping in the model ($|t_{\perp}^{z}|\approx 0.635$ eV) [6], this gives a moderate $U/|t_{\perp}^{z}|\approx 1.83$, reflecting a itinerant electron picture. For these parameters $U_{\perp}\!=\!0$ the leading and subleading solutions of Eq. (7) display $s^{\pm}$-wave and $d_{xy}$-wave symmetries (with $\lambda^{s^{\pm}}/\lambda^{d_{xy}}>1$), in accordance with previous studies [11, 14, 8].
For increasing values of $U_{\perp}$, however, the ratio $\lambda^{s^{\pm}}/\lambda^{d_{xy}}$ decreases and eventually the solution with $d_{xy}$-wave symmetry becomes dominant. This is illustrated in Figure 2(a), which shows a color map of $\log{(\lambda^{s^{\pm}}/\lambda^{d_{xy}})}$ for different values of onsite $U$ and interlayer $U_{\perp}$ interaction strengths. The dashed line represents the curve $\lambda^{s^{\pm}}=\lambda^{d_{xy}}$ and can be interpreted as a “phase boundary” separating the regions where each gap symmetry dominates.
For the reference value of $U\!=\!1.16$ eV, we find that the transition point occurs at a moderate value of $U^{c}_{\perp}\approx 0.2U=0.23$ eV, with the gap functions displaying $s^{\pm}$-wave and $d_{xy}$-wave symmetries for $U_{\perp}<U^{c}_{\perp}$ and $U_{\perp}>U^{c}_{\perp}$ (Figs. 2(b) and (c)), respectively. For larger values of $U$, however, the $s^{\pm}$-wave symmetry is clearly favored even for ratios $U_{\perp}/U$ as large as $50-60\%$.
Figure 3: Color maps of the leading superconducting gap function $\Delta_{\rm max}(\bm{k})$ computed using the two-orbital bilayer pairing vertex but considering only the contributions of two out of the three FS pockets. Panels (a) and (b) shows results excluding the contributions of the $\gamma$ pocket for $U_{\perp}=0$ and $U_{\perp}=0.3U=0.348$ eV, respectively. Panels (c) and (d) show results for $U_{\perp}=0$ excluding the contributions of $\alpha$ and $\beta$ pockets, respectively. In all cases, $U=1.16$ eV.
Some insight as to why the interlayer interaction tends to favor a d-wave symmetry over the $s^{\pm}$ one can be gained by looking more closely on the role of the interorbital contribution to the pairing interaction. Imagine if one could tune the relative contributions of the different FS sheets to the paring gap function defined by Eq. (7). In practice, this can be done by artificially giving different weights to the terms in the sum on the left-hand-side of Eq. (7). For instance, one could exclude (assign zero weight to) the contribution of one of the FS pockets ($\alpha$, $\beta$, or $\gamma$) and see what the resulting symmetry of the leading pairing gap function looks like.
This is done in Fig. 3 for $U=1.16$ eV. Figure 3(a) shows that, for $U_{\perp}=0$, the paring gap symmetry remains $s^{\pm}$-wave even if the contribution of the $\gamma$ pocket is excluded and only pockets $\alpha$ and $\beta$, which are dominated by $3d_{x^{2}-y^{2}}$ orbitals and hence have a more “planar” character, are considered in the calculation. In a sense, this would mimic a “cuprate-like” scenario [16]. In this situation, by increasing $U_{\perp}$ the symmetry of the leading gap function becomes $d_{x^{2}-y^{2}}$ and not $d_{xy}$, as in the case where all pockets are considered (Fig. 2(c)). One the other hand, the pairing symmetry becomes $d_{xy}$ if, instead, the contribution of the $\alpha$ pocket is excluded (Fig. 3(c)). Even though Fig. 3(c) shows results for $U_{\perp}=0$, we have checked that this result holds for a large range of $U_{\perp}$ values. Finally, if the contribution of the $\beta$ pocket is excluded, the gap symmetry becomes $d_{x^{2}-y^{2}}$ (Fig. 3(d)) and not $d_{xy}$.
These results indicate that, when all three pockets are considered in the calculation (Fig. 2), by increasing $U_{\perp}$, the relative contribution of the $\gamma$ pocket (dominated by $3d_{z^{2}}$ orbitals) increases over the (more planar) $\alpha$ pocket contribution, thereby inducing the switching between $s^{\pm}$- and nodal $d_{xy}$-wave paring symmetries. As such, the interplay between $\gamma$ and $\alpha$ pockets in the pairing vertex is a crucial element in explaining the $s^{\pm}\rightarrow d_{xy}$ order switching as $U_{\perp}$ increases.
Figure 4: Enhancement of the pairing vertex function ($\Gamma(\bm{k},\bm{k}^{\prime})$) by interlayer interactions. Panels (a) and (b) show the ratio of $\Gamma(\bm{k},\bm{k}^{\prime})$ calculated with $U_{\perp}=0.348$ eV and $\Gamma^{\mathrm{on}}(\bm{k},\bm{k}^{\prime})$, calculated with $U_{\perp}=0$ in a logarithmic scale for a fixed $\bm{k}$ (black ring) located at (a) $\gamma$ and (b) $\alpha$ Fermi surface pockets, dominated by $d_{3z^{2}-r^{2}}$ and $d_{x^{2}-y^{2}}$ orbitals respectively.
This conclusion is further supported by looking at the differences in the pairing vertex function $\Gamma(\bm{k},\bm{k}^{\prime})$ for $U_{\perp}=0.348$ eV and $U_{\perp}=0$. Figure 4 shows the ratio between $\Gamma(\bm{k},\bm{k}^{\prime})$ (calculated with $U_{\perp}=0.348$ eV) and $\Gamma^{\rm on}(\bm{k},\bm{k}^{\prime})$ (calculated with $U_{\perp}=0$), for a fixed $\bm{k}$ (marked as a black circle) and as a function of $\bm{k}^{\prime}$ along the FS.
For example, as shown in Fig., 4(a), for a $\bm{k}$-point located in the $\gamma$ pocket, increasing $U_{\perp}$ leads to an enhancement of the pairing vertex function ($\Gamma(\bm{k},\bm{k}^{\prime})>\Gamma^{\rm on}(\bm{k},\bm{k}^{\prime})$) along the $\alpha$ pocket and a decrease ($\Gamma(\bm{k},\bm{k}^{\prime})<\Gamma^{\rm on}(\bm{k},\bm{k}^{\prime})$) along the $\beta$ pocket. A similar result occurs when $\bm{k}$ is located at the $\alpha$ pocket (Fig, 4(b)): the pairing vertex function is enhanced along the $\gamma$ pocket and decreases along the $\beta$ pocket.
This result is consistent with our picture that interlayer interactions tend to favor the $d_{xy}$ pairing symmetry over the $s^{\pm}$ by favoring the interorbital pairing between electrons in the $\gamma$ and $\alpha$ pockets, which are dominated by the $d_{3z^{2}-r^{2}}$ and $d_{x^{2}-y^{2}}$ orbital contributions, respectively.
Concluding remarks.— In this work, we investigate the role of interlayer Coulomb interaction terms on the superconducting properties of $\text{La}_{3}\text{Ni}_{2}\text{O}_{7}$ under pressure. Our results indicate that for moderate onsite interactions in the model (up to $U\sim 1.1$ eV), and relatively small interlayer interaction strengths (about $\sim 30\%$ of the onsite term) the leading pairing symmetry becomes of the $d_{xy}$-wave type instead of the $s^{\pm}$-wave predicted by several works in the absence of interlayer interactions [11, 14, 16, 15, 11, 12, 31, 32, 8]. Our calculations also confirm that the $s^{\pm}$ solution is dominant for large onsite interactions ($U\gtrsim 1.3$ eV), which agrees with previous calculations based on $t-J-J_{\perp}$ models [14].
More interestingly, our analysis shows that the superconducting pairing symmetry is strongly dependent on the interplay of all three Fermi surface pockets. In particular, the interlayer interaction tends to favor interorbital pairing between electrons with $d_{3z^{2}-r^{2}}$ and $d_{x^{2}-y^{2}}$ orbital character, underscoring the view that a complete description of the superconducting properties can only be achieved by multi-orbital models [8, 10, 16, 9].
We acknowledge financial support from the São Paulo Research Foundation (FAPESP), Brazil (process numbers 2022/15453-0 and 2023/14902-8) and from CNPq (Grant No. 312622/2023-6).
S1 Matrix random-phase approximation
Since our investigation focuses on exploring the role of multiorbital effects on the magnetic fluctuations of the pressurized bilayer nickelate, we consider the two-orbital $d_{x^{2}-y^{2}}$ and $d_{z^{2}-r^{2}}$ low-energy model [6], as represented in Fig. 1(a) of the main text. The non-interacting Hamiltonian reads
$\begin{split}H_{0}&=\sum_{\bm{k}\sigma sS}(T_{\bm{k}}^{s}-\mu)c_{\bm{k}sS\sigma}^{\dagger}c_{\bm{k}sS\sigma}\\ &+\sum_{\bm{k}\sigma s}t_{\perp}^{s}(c_{\bm{k}sA\sigma}^{\dagger}c_{\bm{k}sB\sigma}+\text{h.c.})\\ &+\sum_{\bm{k}\sigma(s\neq p)S}V_{\bm{k}}c_{\bm{k}sS\sigma}^{\dagger}c_{\bm{k}pS\sigma}\\ &+\sum_{\bm{k}\sigma(s\neq p)}V_{\bm{k}}^{\prime}(c_{\bm{k}sA\sigma}^{\dagger}c_{\bm{k}pB\sigma}+\text{h.c.}),\end{split}$ (S1)
where $c^{\dagger}_{\bm{k}sS\sigma}$ creates an electron with spin $\sigma$, in the orbital $s=\{x,z\}$ corresponding to $\{d_{x^{2}-y^{2}},d_{3z^{2}-r^{2}}\}$, sublattice $S=\{A,B\}$, and momentum $\bm{k}$, $\mu$ is the chemical potential, and h.c. shortcuts Hermitian conjugate. This model considers up to second-neighbor hoppings such that
$\begin{split}T_{\bm{k}}^{s}&=2t_{1}^{s}(\cos k_{x}+\cos k_{y})+4t_{2}^{s}\cos k_{x}\cos k_{y}+\epsilon^{s}\\ V_{\bm{k}}&=2t_{3}^{xz}(\cos k_{x}-\cos k_{y}),\\ V_{\bm{k}}^{\prime}&=2t_{4}^{xz}(\cos k_{x}-\cos k_{y}).\end{split}$ (S2)
The hopping parameters are given in Ref. [6]. Fig. 1(c) of the main text shows the band structure with its respective orbital composition. There is an interplay between the two orbitals $d_{x^{2}-y^{2}}$ (red) and $d_{3z^{2}-r^{2}}$ (blue) along the path $\bm{\Gamma}-\bm{X}-\bm{M}$. Also, the $\gamma$ pocket is dominated by the $d_{3z^{2}-r^{2}}$ orbital and contributed by a nearly flat hole pocket centered at the $\bm{M}$ point.
We consider interactions through a degenerate Hubbard model including interlayer interactions. Within this model, the electrons are not constrained to be localized and the specific values of the interaction strengths are not restricted to the limit $U/t\gg 1$ such as the $t-J$-like models considered in the previous literature. In fact, it has been estimated using the random-phase approximation within mean-field theory that $U\approx 1.16$ eV to match $T_{c}=80$ K [liu\$s^ifmmodepmelsetextpmfi\$WavePairingDestructive2023], while the largest hopping is $t_{\perp}^{z}\approx-0.635$ eV [6] such that $U/t_{\perp}^{z}\approx-1.83$, which is a moderate value. The interaction Hamiltonian for the two-orbital itinerant electron model is given in the main text. Since we are interested in the magnetic fluctuations caused by interacting electrons, we focus our study on the bare and interacting susceptibilities. Having in mind that the present non-interacting model [Eq. (S1)] is independent of the spin directions, the bare susceptibility is the same for charge or spin channels and can be defined as [19]
[…]
where $i\nu_{m}$ denotes bosonic frequencies, $\tau$ imaginary time, $T$ temperature, and $T_{\tau}$ time-ordering. We compute the bare susceptibility in the Green’s function form performing the Matsubara summation numerically using Ozaki’s method [33]. The formalism reads
$\begin{split}&[\hat{\chi}]^{pAqB}_{sCtD}(\bm{q},i\nu_{m})=-\frac{T}{N}\sum_{\bm{k}i\omega_{n}}G_{sCpA}(\bm{k},i\omega_{n})G_{qBtD}(\bm{k}+\bm{q},i\omega_{n}+i\nu_{m}),\end{split}$ (S4)
where
$G_{sCpA}(\bm{k},i\omega_{n})=\sum_{\nu_{\bm{k}}}\frac{\psi^{sC}_{\nu_{\bm{k}}}(\bm{k})\psi^{pA*}_{\nu_{\bm{k}}}(\bm{k})}{i\omega_{n}-E_{\nu_{\bm{k}}}(\bm{k})},$ (S5)
and $\psi^{sC}_{\nu_{\bm{k}}}(\bm{k})$ are non-interacting electron eigenfunctions related to the $\nu_{\bm{k}}$-th $E_{\nu_{\bm{k}}}(\bm{k})$ energy band of Eq. (S1). We adopt $T=0.02$ eV.
Within the mRPA, adding the onsite and interlayer interactions reduces to performing a matrix operation such that the spin and charge susceptibilities become
$\hat{\chi}_{s}(\bm{q},i\nu_{m})$ $=\hat{\chi}(\bm{q},i\nu_{m})[\hat{1}-\hat{U}_{s}\hat{\chi}(\bm{q},i\nu_{m})]^{-1},$ (S6)
$\hat{\chi}_{c}(\bm{q},i\nu_{m})$ $=\hat{\chi}(\bm{q},i\nu_{m})[\hat{1}+\hat{U}_{c}\hat{\chi}(\bm{q},i\nu_{m})]^{-1}.$ (S7)
S2 Pairing symmetry of individual pockets
As discussed in Sec. Superconducting symmetries of the main text, the role of considering a two-orbital model for the superconductivity of bilayer nickelates can be tested by solving the eigenvalue Eq. (7) for the desired pockets only. In the argumentation, we note that we have checked that no individual pocket shows the desired $s^{\pm}$- and $d_{xy}$-wave interplay as the two-orbital model. The calculations are shown in Fig. S1. Panels (a) and (b) show no change in the $d_{xy}$ symmetry of the $\beta$ pocket as the interlayer coupling varies from $U_{\perp}/U=0$ (a) to $U_{\perp}/U=0.3$ (b). In fact, the interlayer interactions do not change the symmetry in any three cases explored. As shown in panel (c), a $d_{x^{2}-y^{2}}$-wave dominates pairing of individual $\gamma$ pocket, and a $g$-wave symmetry is obtained for the $\alpha$ pocket alone (d).
Figure S1: Same as Fig. 3 of the main text but for individual pockets $\beta$ (a,b), $\gamma$ (c), and $\alpha$ (d).
Figure S2: Cuts of the phase diagram shown in Fig. 2 in the main text.